# Adiabatic modulation of driving protocols in periodically driven quantum systems

### v1

@@ Abstract
We consider a periodically driven system where the high-frequency driving protocol consists of a sequence of potentials switched on and off at different instants within a period. We explore the possibility of introducing an adiabatic modulation of the driving protocol by considering a slow evolution of the instants when the sequence of potentials is switched on/off. We examine how this influences the long-term dynamics of periodically driven quantum systems. By assuming that the slow and fast timescales in the problem can be decoupled, we derive the stroboscopic (effective) Hamiltonian for a four-step driving sequence up to the first order in perturbation theory. We then apply this approach to a rigid rotor, where the adiabatic modulation of the driving protocol is chosen to produce an evolving emergent magnetic field that interacts with the rotor’s spin. We study the emergence of diabolical points and diabolical loci in the parameter space of the effective Hamiltonian. Further, we study the topological properties of the maps of the adiabatic paths in the parameter space to the eigenspace of the effective Hamiltonian. In effect, we obtain a technique to tune the topological properties of the eigenstates by selecting various adiabatic evolution of the driving protocol characterized by different paths in the parameter space. This technique can be applied to any periodic driving protocol to achieve desirable topological effects.
@@ I Introduction
A class of systems involving the coupling of a spin with a magnetic field is of great interest as they demonstrate non-trivial topological properties [21, 22, 23, 24, 25]. The eigenstates of such systems generally acquire a geometrical phase [26, 27, 28, 29, 30] when slowly transported around a closed path by varying the parameters $\vec{R}$ in its Hamiltonian ^$\widehat{\mathcal{H}}(\vec{R})$. Further, degeneracies in such systems with at least two parameters also exhibit diabolical points with double-cone-like structures at such points in the band diagram [31, 32, 33]. In the traditional studies of adiabatic evolution [34, 35], the parameter space usually comprises of the constants, which are the strengths of the different operators appearing in the Hamiltonian. It is not usually possible to increase the number of parameters without introducing new operators which thereby limits the complexity of the dynamics. We suggest a novel method of synthesizing a Hamiltonian where the number of parameters can be increased without requiring any new basic operators as the building blocks for the Hamiltonian.
@@
In this paper, we first consider a high-frequency driving of a spin-$1/2$ rotor with a protocol consisting of a four-step sequence of potentials switched on and off at different instants within a period. We further invoke a much slower time scale for an adiabatic modulation of the driving protocol itself through a slow variation of the instants when the sequence of potentials is switched on/off. The choice of operators in the driving protocol produces an evolving emergent magnetic field that interacts with the rotor’s spin. We then look at adiabatic evolution in the space of the parameters of the driving protocol, which effectively maps to the evolution of the synthesized magnetic field. We study the emergence of diabolical points and topological properties of adiabatic paths. The geometric phase picked up by the eigenstates during the adiabatic evolution and its variation for different forms of adiabatic protocols is subsequently studied. In effect, we obtain a technique to tune the winding of the eigenstates and the accumulated geometric phase by choosing various adiabatic evolution of the driving protocol characterized by diverse paths in the parameter space.
@@ II Formalism
$\widehat{\mathcal{U}}(t_{i},t_{f})=e^{-i\widehat{\mathcal{K}}(t_{f})}e^{-i\widehat{\mathcal{H}}_{\text{eff}}(t_{f}-t_{i})}e^{i\widehat{\mathcal{K}}(t_{i})}.$ (2)
Here, $\widehat{\mathcal{K}}(t)$ is a time-periodic operator with period $T$ known as the micro-motion kick operator, and ^$\widehat{\mathcal{H}}_{\text{eff}}$ is a time-independent effective Hamiltonian. The expressions for the kick operator and the effective Hamiltonian can be obtained using a perturbative expansion [19, 18] in the limit of high driving frequency $\omega=\frac{2\pi}{T}>>\omega_{0}$, where $\hbar\omega_{0}$ is the typical energy spacing of the base Hamiltonian ^$\widehat{\mathcal{H}}_{0}$.
@@
We follow the formalism in Appendix C of [19] where the the perturbative terms ^$\widehat{\mathcal{H}}_{m}$ of the effective Hamiltonian ^$\widehat{\mathcal{H}}_{\text{eff}}$ are computed using the Fourier components $V^{(j)}$ of the driving potential defined through
@@
^$\widehat{\mathcal{H}}_{\text{eff}}=\widehat{\mathcal{H}}_{0}+\frac{i\pi}{8\omega}\sum_{\begin{subarray}{c}r,s=1\\ s>r\end{subarray}}^{4}\mathscr{P}_{rs}\left(\alpha,\beta\right)\left[\widehat{V}_{r},\widehat{V}_{s}\right]+\mathcal{O}\left(\omega^{-2}\right),$ (6)
@@
The effective Hamiltonian ^$\widehat{\mathcal{H}}_{\text{eff}}$ is obtained by averaging over the high-frequency variation of the potential. The fast-driving protocol can be morphed by changing $(\alpha,\beta)$ through the changes in the values of the polynomials $\mathscr{P}_{rs}(\alpha,\beta)$. We thus consider, another time scale $>>\omega^{-1}$ over which $(\alpha,\beta)$ evolve. This results in an adiabatic evolution of ^$\widehat{\mathcal{H}}_{\text{eff}}(\vec{R}(\tau))$, where $\vec{R}(\tau)=(\alpha(\tau),\beta(\tau))\in[0,1]\times[0,1]$ evolves smoothly and slowly in the parameter space. We use $\tau$ to indicate that we looking at adiabatic variations different from the fast stroboscopic time scale $t$ varying over $T$. This is akin to Berry’s adiabatic evolution [27] of a Hamiltonian in the parameter space, except that our parameters are now the characteristics of a fast-driving protocol. The adiabatic evolution $\vec{R}(\tau)=(\alpha(\tau),\beta(\tau))$ is a path in the 2D parameter space $[0,1]\times[0,1]\subset{\mathbb{R}}^{2}$. We parameterize this path by a continuous vector-valued map
@@
We are interested in the possibility that the eigenstates of ^$\widehat{\mathcal{H}}_{\text{eff}}(\vec{R}(\tau))$ picks up nontrivial geometric phase under such adiabatic evolution.
II.1 Driven rigid rotor
@@
Adiabatic evolution of quantum Hamiltonian systems [39, 40] in the parameter space is comprehensively investigated for undriven Hamiltonians, where the adiabatic parameters are restricted to the parameters of the stationary Hamiltonian in the manner of Berry [27]. We are interested in driven systems and aim to use the parameters of the driving protocol to induce adiabatic evolution of an effective Hamiltonian. One of the simplest systems demonstrating anholonomy is the two-level spin-$1/2$ system, subjected to an external magnetic field, for which the ground state picks up a non-zero Berry phase on adiabatically varying the direction of the magnetic field. This motivates us to engineer a driving protocol that would give us an effective Hamiltonian like the anholonomy Hamiltonian of the two-state system. That is, we desire that the engineered Hamiltonian is of the form $\mathbf{J}\cdot\mathbf{B}$, where $\mathbf{B}$ is an effective magnetic field.
We choose our undriven system to be a rigid rotor with spin $\mathbf{J}$, such that ^$\widehat{\mathcal{H}}_{0}=\frac{J^{2}}{2I}$. We drive this system with a protocol $\Pi\left(\alpha,\beta,c_{i}\right)$ with the potentials given by:
$\widehat{V}_{1}=c_{1}J_{x}-c_{2}J_{y}\ ,\widehat{V}_{2}=c_{2}J_{y}+c_{3}J_{z},$
$\widehat{V}_{3}=c_{4}J_{z}-c_{1}J_{x}\ ,\widehat{V}_{4}=-(c_{3}+c_{4})J_{z},$ (11)
@@
^$\widehat{\mathcal{H}}_{\text{eff}}$ $=\frac{J^{2}}{2I}-\frac{\pi}{8\omega}\mathbf{J}\cdot\mathbf{B}\quad\text{and}\quad\widehat{\mathcal{K}}\left(0\right)=-\frac{\pi}{4\omega}\mathbf{J}\cdot\mathbf{B^{\prime}},$ (12)
@@ III.1 The Energy Spectrum
The eigenvalues of the effective Hamiltonian ^$\widehat{\cal{H}}_{\text{eff}}$ as a function of the parameters $(\alpha,\beta)$ are given by $E(\alpha,\beta)=\pm{|\mathbf{B}|}$. A choice of parameters $(\alpha,\beta)$, thus allows us to tune the spectrum by causing level repulsion or reduction of the energy gap. The gapless degenerate points are particularly important since they are topologically protected.
@@ III.2 Adiabatic Paths
Thus, any back-and-forth evolution of $(\alpha,\beta)$ on each of the four segments of the square boundary and also the leading diagonal will not lead to any phase picked up by $|\chi\rangle$ since they correspond to fixed directions of $\hat{n}$. Further, any closed path in the parameter space that crosses any of these five segments more than once will correspond to closed paths that cross itself on $S^{2}$. The crossing point on the sphere corresponds to the two points where $\hat{n}$ is the same owing to $(\alpha,\beta)$ lying on any of these five segments. Thus, closed paths in the parameter space that cross these segments multiple times at different places will lead to a winding of $|\chi\rangle$ on $S^{2}$.
@@
This path crosses one of the five segments, namely the leading diagonal at the points $p_{1}$ and $p_{2}$. Since $p_{1}$ and $p_{2}$ correspond to the same eigenstate $p$ on $S^{2}$, the path on $S^{2}$ winds twice as $\vec{R}\left(\tau\right)$ winds once in the parameter space as shown in Fig. 5(e). In general, depending on the number of times a simple closed path on the parameter space intersects the dashed segments, we could have multiple winding of the eigenstate on $S^{2}$. The geometric phase picked would depend on the orientation of the path on $S^{2}$ as it evolves. In principle, the path in the parameter space can be constructed so that there is no net geometric phase picked up under such an evolution.
We next consider an open path $p_{1}\rightarrow p_{2}\rightarrow p_{3}\rightarrow p_{4}$ on the parameter space that corresponds to an adiabatic evolution of the state with multiple winding, as shown in Fig. 5$\left(c\right)$ given by
@@
This path symmetrically intersects two of the dashed segments at $p_{1},p_{4}$ and $p_{2},p_{3}$ respectively. This would result in the eigenstate winding twice on $S^{2}$ as seen in Fig. 5(f). Thus, even for an open path in the parameter space, we can have the eigenstate cyclically trace a closed path on $S^{2}$ and pick up a phase in the process. The points $p_{1}$ and $p_{4}$ lie on a contour of constant magnitude of the magnetic field $\|\mathbf{B}\|=1.24$. This implies that both the magnitude and the direction of the synthesized magnetic field return to the initial value at the end of the evolution.
@@ IV.1 Winding number
We have seen that for the protocol $(c_{1},c_{2},c_{3},c_{4})=(1,1,1,1)$, a closed trajectory in the parameter space that crosses the leading diagonal (the line $\alpha=\beta$) or any of the boundary lines shall cause the eigenstate of the effective Hamiltonian to loop more than one on the Bloch sphere. The winding number for the eigenstate is then determined by the number of times the trajectory crosses these lines. We have further established in the last section that the energy cost of a trajectory is path-dependent. Since $\Delta E_{\text{s}low}$ is a smooth function of the path $\vec{R}(\tau)$, we claim that for two trajectories that are almost the same, the energy cost shall also be almost the same with a difference that shall be proportional to the tiny difference of the path lengths. We demonstrate this using the trajectory in Fig. 5(a)
Figure 6: Two closed paths $MNOPQR$ and $PQR$ in the parameter space which differ marginally. However, for path $MNOPQR$ the winding number of the eigenstate on $S^{2}$ is 2, while for the path $PQR$ it is 1.
In Fig. 6, the two closed paths $MNOPQR$ and $PQR$ in the parameter space differ slightly. Thereby the energy costs for evolution along these paths will almost be the same. However, the path $MNOPQR$ crosses the diagonal line at two points, and thus, for this path, the eigenstate shall have a winding number of two on $S^{2}$. This demonstrates that it is possible to generate higher winding numbers by smooth variation of the path in parameter space, which does not cost much energy. The discrete jump of the winding number is reminiscent of a phase transition [34, 36, 42] and may have topological significance.

### v2

@@ Abstract
We consider a periodically driven system where the high-frequency driving protocol consists of a sequence of potentials switched on and off at different instants within a period. We explore the possibility of introducing an adiabatic modulation of the driving protocol by considering a slow evolution of the instants when the sequence of potentials is switched on/off. We examine how this influences the long-term dynamics of periodically driven quantum systems. By assuming that the slow and fast timescales in the problem can be decoupled, we derive the stroboscopic (effective) Hamiltonian for a four-step driving sequence up to the first order in perturbation theory. We then apply this approach to a spin-$1/2$ system, where the adiabatic modulation of the driving protocol is chosen to produce an evolving emergent magnetic field that interacts with the spin. We study the emergence of diabolical points and diabolical loci in the parameter space of the effective Hamiltonian. Further, we study the topological properties of the maps of the adiabatic paths in the parameter space to the eigenspace of the effective Hamiltonian. In effect, we obtain a technique for tuning the topological properties of the eigenstates by selecting different adiabatic evolutions of the driving protocol, characterized by distinct paths in parameter space. This technique can be applied to any periodic driving protocol to achieve desirable topological effects.
@@ I Introduction
A class of systems involving the coupling of a spin with a magnetic field is of great interest as they demonstrate non-trivial topological properties [21, 22, 23, 24, 25]. The eigenstates of such systems generally acquire a geometrical phase [26, 27, 28, 29, 30] when slowly transported around a closed path by varying the parameters $\vec{R}$ in its Hamiltonian ^$\widehat{\mathcal{H}}(\vec{R})$. Further, degeneracies in such systems with at least two parameters also exhibit diabolical points with double-cone-like structures at such points in the band diagram [31, 32, 33]. In traditional studies of adiabatic evolution [34, 35], the parameter space consists of the constants representing the strengths of different operators appearing in the Hamiltonian. It is usually not possible to increase the number of parameters without introducing new operators, thereby limiting the complexity of the dynamics. We suggest a novel method of synthesizing a Hamiltonian where the number of parameters can be increased without requiring any new basic operators as the building blocks for the Hamiltonian.
@@
In this paper, we first consider a high-frequency driving of a spin-$1/2$ system with a protocol consisting of a four-step sequence of potentials switched on and off at different instants within a period. We further invoke a much slower time scale for an adiabatic modulation of the driving protocol itself through a slow variation of the instants when the sequence of potentials is switched on/off. The choice of operators in the driving protocol produces an evolving emergent magnetic field that interacts with the spin. We then look at adiabatic evolution in the space of the parameters of the driving protocol, which effectively maps to the evolution of the synthesized magnetic field. We study the emergence of diabolical points and topological properties of adiabatic paths. The geometric phase picked up by the eigenstates during the adiabatic evolution and its variation for different forms of adiabatic protocols is subsequently studied. In effect, we obtain a technique to tune the winding of the eigenstates and the accumulated geometric phase by choosing various adiabatic evolution of the driving protocol characterized by diverse paths in the parameter space.
@@ II Formalism
$\widehat{\mathcal{U}}(t_{i},t_{f})=e^{-\mathrm{i}\widehat{\mathcal{K}}(t_{f})}e^{-\mathrm{i}\widehat{\mathcal{H}}_{\text{eff}}(t_{f}-t_{i})}e^{\mathrm{i}\widehat{\mathcal{K}}(t_{i})}.$ (2)
Here, $\widehat{\mathcal{K}}(t)$ is a time-periodic operator with period $T$ known as the micro-motion kick operator, and ^$\widehat{\mathcal{H}}_{\text{eff}}$ is a time-independent effective Hamiltonian. The expressions for the kick operator and the effective Hamiltonian can be obtained using a perturbative expansion [19, 18] in the limit of high driving frequency $\omega=\frac{2\pi}{T}>>\omega_{0}$, where $\hbar\omega_{0}$ is the typical energy scale of the base Hamiltonian ^$\widehat{\mathcal{H}}_{0}$.
@@
We follow the formalism in Appendix C of [19] where the perturbative terms ^$\widehat{\mathcal{H}}_{m}$ of the effective Hamiltonian ^$\widehat{\mathcal{H}}_{\text{eff}}$ are computed using the Fourier components $V^{(j)}$ of the driving potential defined through
@@
^$\widehat{\mathcal{H}}_{\text{eff}}=\widehat{\mathcal{H}}_{0}+\frac{\mathrm{i}\pi}{8\omega}\sum_{\begin{subarray}{c}r,s=1\\ s>r\end{subarray}}^{4}\mathscr{P}_{rs}\left(\alpha,\beta\right)\left[\widehat{V}_{r},\widehat{V}_{s}\right]+\mathcal{O}\left(\omega^{-2}\right),$ (6)
@@
The effective Hamiltonian ^$\widehat{\mathcal{H}}_{\text{eff}}$ is obtained by averaging over the high-frequency variation of the potential. The fast driving protocol can be modified by changing $(\alpha,\beta)$ through changes in the values of polynomials $\mathscr{P}_{rs}(\alpha,\beta)$. We therefore consider another time scale $\tau>>\omega^{-1}$ over which $(\alpha,\beta)$ evolve. This results in an adiabatic evolution of ^$\widehat{\mathcal{H}}_{\text{eff}}(\vec{R}(\tau))$, where $\vec{R}(\tau)=(\alpha(\tau),\beta(\tau))\in[0,1]\times[0,1]$ evolves smoothly and slowly in the parameter space. We use $\tau$ to indicate that we looking at adiabatic variations different from the fast stroboscopic time scale $t$ varying over $T$. This is akin to Berry’s adiabatic evolution [27] of a Hamiltonian in the parameter space, except that our parameters are now the characteristics of a fast-driving protocol. The adiabatic evolution $\vec{R}(\tau)=(\alpha(\tau),\beta(\tau))$ is a path in the 2D parameter space $[0,1]\times[0,1]\subset{\mathbb{R}}^{2}$. We parameterize this path by a continuous vector-valued map
@@
We are interested in the possibility that the eigenstates of ^$\widehat{\mathcal{H}}_{\text{eff}}(\vec{R}(\tau))$ pick up a nontrivial geometric phase under such adiabatic evolution.
II.1 Driven spin $1/2$ system
@@
Adiabatic evolution of quantum Hamiltonian systems [39, 40] in the parameter space is comprehensively investigated for undriven Hamiltonians, where the adiabatic parameters are restricted to the parameters of the stationary Hamiltonian in the manner of Berry [27]. We are interested in driven systems and aim to use the parameters of the driving protocol to induce adiabatic evolution of an effective Hamiltonian. One of the simplest systems demonstrating anholonomy is the two-level spin-$1/2$ system, subjected to an external magnetic field, for which the ground state picks up a non-zero Berry phase on adiabatically varying the direction of the magnetic field. This motivates us to engineer a driving protocol that would give us an effective Hamiltonian like the anholonomy Hamiltonian of the two-state system. That is, we desire that the engineered Hamiltonian is of the form $\mathbf{S}\cdot\mathbf{B}$, where $\mathbf{B}$ is an effective magnetic field and $\mathbf{S}=\left(S_{x},S_{y},S_{z}\right)$ is the spin angular momentum operator.
We choose our undriven system to be a two-level system of a spin-$1/2(s=1/2)$, particle, such that ^$\widehat{\mathcal{H}}_{0}=\frac{3}{4}\frac{\mathds{1}_{\text{2x2}}}{2I}$ in the $S_{z}$ basis. The dimensionful number $I$ sets the scale of energy for the undriven system. We drive this system with a protocol $\Pi\left(\alpha,\beta,c_{i}\right)$ with the potentials given by:
$\widehat{V}_{1}=c_{1}S_{x}-c_{2}S_{y}\ ,\widehat{V}_{2}=c_{2}S_{y}+c_{3}S_{z},$
$\widehat{V}_{3}=c_{4}S_{z}-c_{1}S_{x}\ ,\widehat{V}_{4}=-(c_{3}+c_{4})S_{z},$ (11)
@@
^$\widehat{\mathcal{H}}_{\text{eff}}$ $=\frac{S^{2}}{2I}-\frac{\pi}{8\omega}\mathbf{S}\cdot\mathbf{B}\quad\text{and}\quad\widehat{\mathcal{K}}\left(0\right)=-\frac{\pi}{4\omega}\mathbf{S}\cdot\mathbf{B^{\prime}},$ (12)
@@
We note that with our choice of ^$\widehat{\mathcal{H}}_{0}$ the perturbative corrections to the energy at and above $\mathcal{O}\left(\omega^{-2}\right)$ vanish.
@@ III.1 The Energy Spectrum
The eigenvalues of the effective Hamiltonian ^$\widehat{\cal{H}}_{\text{eff}}$ as a function of the parameters $(\alpha,\beta)$ are given by $E\left(\alpha,\beta\right)=\frac{3}{8I}\pm\frac{\pi}{8\omega}\lvert\mathbf{B}(\alpha,\beta)\rvert$. Upto an overall shift of energy by $\frac{3}{8I}$ and scaling by $\frac{\pi}{8\omega}$ the spectrum is determined by ±$\pm{\lvert\mathbf{B(\alpha,\beta)}\rvert}$. We use units $\hbar=c=1$ in our analysis, such that $E\left(\alpha,\beta\right)$ and $\lvert\mathbf{B\left(\alpha,\beta\right)}\rvert$ are dimensionless quantities with all energies in units of $I^{-1}$. A choice of parameters $(\alpha,\beta)$, thus allows us to tune the spectrum by causing level repulsion or reduction of the energy gap. The gapless degenerate points are particularly important since they are topologically protected.
@@ III.2 Adiabatic Paths
Thus, any back-and-forth evolution of $(\alpha,\beta)$ on each of the four segments of the square boundary and also the leading diagonal will not lead to any phase picked up by $|\chi\rangle$ since they correspond to fixed directions of $\hat{n}$. Further, any closed path in the parameter space that crosses any of these five segments more than once will correspond to closed paths that cross itself on $S^{2}$. The crossing point on the sphere corresponds to the two points where $\hat{n}$ is the same owing to $(\alpha,\beta)$ lying on any of these five segments. Thus, closed paths in the parameter space that cross these segments multiple times at different places will lead to one or many crossings of |$\ket{\chi}$ in $S^{2}$. Noting that a closed trajectory in the parameter space is mapped to a closed path on the Bloch sphere, the number of crossings is an integer which characterizes the map from $S^{1}\to S^{1}$ since $\pi_{1}\left(S^{1}\right)=\mathbb{Z}$.
@@
This path crosses one of the five segments, namely the leading diagonal at the points $p_{1}$ and $p_{2}$. Since $p_{1}$ and $p_{2}$ correspond to the same eigenstate $p$ on $S^{2}$, the path on $S^{2}$ winds twice as $\vec{R}\left(\tau\right)$ winds once in the parameter space as shown in Fig. 5(e). In general, depending on the number of times a simple closed path on the parameter space intersects the dashed segments, we could have multiple crossings of the eigenstate on $S^{2}$. The geometric phase picked would depend on the orientation of the path on $S^{2}$ as it evolves. In principle, the path in the parameter space can be constructed so that there is no net geometric phase picked up under such an evolution.
We next consider an open path $p_{1}\rightarrow p_{2}\rightarrow p_{3}\rightarrow p_{4}$ on the parameter space that corresponds to an adiabatic evolution of the state with multiple windings, as shown in Fig. 5$\left(c\right)$ given by
@@
This path symmetrically intersects two of the dashed segments at $p_{1},p_{4}$ and $p_{2},p_{3}$ respectively. This would result in the eigenstate winding twice on $S^{2}$ as seen in Fig. 5(f). Thus, even for an open path in the parameter space, we can have the eigenstate cyclically trace a closed path on $S^{2}$ and pick up a phase in the process. The points $p_{1}$ and $p_{4}$ lie on a contour of constant magnitude of the magnetic field $\lvert\mathbf{B}\rvert=1.24$. This implies that both the magnitude and the direction of the synthesized magnetic field return to the initial value at the end of the evolution.
@@ IV.1 Winding number
We have seen that for the protocol $(c_{1},c_{2},c_{3},c_{4})=(1,1,1,1)$, a closed trajectory in the parameter space that crosses the leading diagonal (the line $\alpha=\beta$) or any of the boundary lines shall cause the eigenstate of the effective Hamiltonian to loop more than one on the Bloch sphere. The winding number for the eigenstate is then determined by the number of times the trajectory crosses these lines. We have further established in the last section that the energy cost of a trajectory is path-dependent. Since this is a smooth function of the path $\vec{R}(\tau)$, we claim that for two trajectories that are almost the same, the energy cost shall also be almost the same with a difference that shall be proportional to the tiny difference of the path lengths. We demonstrate this using the trajectory in Fig. 5(a)
Figure 6: Two closed paths $MNOPQR$ and $PQR$ in the parameter space which differ marginally. However, for path $MNOPQR$ the winding number is 2 and the eigenstate will cross itself twice, while for the path $PQR$ it is 1.
In Fig. 6, the two closed paths $MNOPQR$ and $PQR$ in the parameter space differ slightly. Thereby the energy costs for evolution along these paths will almost be the same. However, the path $MNOPQR$ crosses the diagonal line at two points, and thus, for this path, the eigenstate shall have a winding number of two. This demonstrates that it is possible to generate higher winding numbers by smooth variation of the path in parameter space, which does not cost much energy. An integer jump in the winding number is intrinsically related to the degeneracy structure of the Hamiltonian in the parameter space. The system chosen by us is devoid of sufficient complexity for physical quantities to undergo a real phase transition with topological significance [34, 36, 42]. The model is aimed simply to demonstrate the possibility of utilizing temporal partitioning as a means to synthesize a non-trivial degeneracy structure in the parameter space. The topological quantity namely the winding number just gives us the number of times the synthesized magnetic field $\vec{B}$ returns back to its original orientation while we go around only once in the parameter space could hold physical significance.